I can’t believe that after all these problems you are still making things! It’s great for
me, because I get to have you do calculations on them. This time, you are making a
solenoid. Naturally, you use a toilet paper tube.3 The tube has a diameter of 4.00 cm
and a height of 11.2 cm. You wrap it with copper wire having a cross-sectional area
of 1.02 × 10−6 m2
(20-gauge; note that the resistivity of copper is 1.70 × 10−8 Ωm ).
Your resulting solenoid has 98 loops. Determine (a) the current required to produce a
magnetic field of 6.00 mT in the center of the solenoid, and (b) the power required to
deliver this magnetic field.
1
Expert's answer
2017-03-28T05:39:06-0400
a) A field of a solenoid is B = mm0*In or I = B/(mm0n) = 5.5A

b) The power of solenoid P = I^2*n*3.14*d*res./S = 6.2 W